You are in the process of deciding the optimal order quantity of shampoo packs for a hotel. The supplier charges you $59 per order and your stockroom costs approximately $0.67 per shampoo pack per year to store the product. Based on prior research, you were able to determine that the hotel goes through approximately 90 shampoo packs per day. They hold 0 safety stock. The supplier is willing to give the following quantity discounts for the product. Using the order quantity that will minimize total cost, what is the total annual cost?

Quantity

Price

1 – 1499

$2.25

1500 – 3999

$2.10

4000 – 7999

$2.05

8000+

$2.03
Answer #1

Given that, Daily demand = 90 Units

Yearly demand = 365 x 90 = 32850 Units

Ordering cost = $59 per Order

Holding cost = $0.67 per unit per year

We know that,

Economic order quantity = EOQ = Sq. Root [(2 x Annual demand x Ordering cost) / Holding cost]

= Sq. Root [(2 x 32850 x 59) / 0.67]

= 2405 Units

Number of orders per year = Annual demand / EOQ = 32850 / 2405 = 13.65 Orders

Annual ordering cost = Number of orders per year x Cost per Order

= 13.65 x 59 = $805.78

In EOQ,

Annual holding cost = Annual Ordering cost = $805.78

Cost per Item = $2.10

Annual cost of purchase = 32850 x 2.10 = $68985

Total cost Annually = Annual cost of purchase + Annual Ordering cost + Annual holding cost

= 68985 + 805.78 + 805.78

= $70596.56